Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
SPB
SQS
XO
ToBeMigrated2XWiki
ToBeMigrated2GitlabPages
ToBeMigratedByGroup
Migrated2Zulip
Archived2PDF
not2MigrateOrArchive
SCS
SXP Migrated2Zulip
Photon Commissioning Team
XRO
DOC
HED Migrated2Zulip
WP74
MID
PSPO_XO
XO_Operation
SQS Migrated2Zulip
FXE Migrated2Zulip
DET
Sample Environment
SPB Migrated2Zulip
OPERATION-2025
OPERATION-2024
OPERATION-2023
OPERATION-2022
OPERATION-2021
OPERATION-2020
OPERATION-2019
SASE1-2018
SASE1-2017
Experiment data
OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn
Not logged in
List
|
Message ID:
113
Entry time:
26 May 2024, 23:23
Author:
Romain Letrun
Group:
SPB
Subject:
Shift summary
X-ray delivery
6.0 keV, up to 202 pulses at 564 kHz, 1.8-1.9 mJ.
Stable delivery except for downtime at start of shift (until 7:15) and two more hours in the morning.
Optical laser delivery
770 nm, up to 202 pulses at 564 kHz.
Stable delivery.
Achievements/Observations
Collected time-resolved diffraction from gold and gold-platinum nanorods at multiple laser fluences.
Issues
Vacuum spike in EHU_PSLIT section.
DAMNIT backend process stuck. A couple of Maxwell nodes were unaccessible. Solved by DOC and DA OCD.
Issue with two AGIPD modules noticed around 21:20. Could not be recovered before end of the shift. To be further investigated tomorrow morning.
ELOG V3.1.4-7c3fd00